Solve and find particular solution of the first order non-linear differential equation of Bernoulli’s type with given condition.

 


1
Expert's answer
2021-02-23T11:35:19-0500

Solution:


the first-order non-linear differential equation of Bernoulli’s type is


"\dfrac{dy}{dx}+P(x)y=Q(x)y^n."


where p(x) and q(x) are continuous functions on the interval we’re working on and n is a real number. Differential equations in this form are called Bernoulli Equations.


Example:


Find the solution of the differential equation "4\nx\ny\ny\n\u2032\n=\ny\n^2\n+\nx\n^2", satisfying the initial condition y(1)=1.


Solution:


First, we should check whether this differential equation is a Bernoulli equation:




"4xyy'=y^2+x^2,\;\rightarrow\;\dfrac{4xyy'}{4xy}-\dfrac{y^2}{4xy}=\dfrac{x^2}{4xy},\;\rightarrow\;y'-\dfrac{y}{4x}=\dfrac{x}{4y}"

As it can be seen, we have a Bernoulli equation with the parameter m=−1. Hence, we can make the substitution "z=y^{1-m}=y^2." The derivative of the function is "z'=2yy'." Next, we multiply both sides of the differential equation by 2y:



"2yy'-\dfrac{2y^2}{4x}=\dfrac{2xy}{4y}, \;\rightarrow\;2yy'-\dfrac{y^2}{2x}=\dfrac{x}{2}."


By replacing y with z, we can convert the Bernoulli equation into the linear differential equation:

"z'=\dfrac{z}{2x}=\dfrac{x}{2}."

Calculate the integrating factor:



"u(x)=e^{\int(-\frac{1}{2x})dx}=e^{{-\frac{1}{2}\int-\frac{dx}x}}=e^{-\frac{1}{2}ln|x|}=\dfrac{1}{\sqrt{|x|}}"

Let`s choose the function "u(x)=\dfrac{1}{\sqrt{x}}" and make sure that the left side of the equation becomes the derivative of the product z(x) u(x) after multiplying by u(x):


"(z'-\dfrac{z}{2x})u(x)=z'\;\cdot\dfrac{1}{\sqrt{x}}-\dfrac{z}{2x}\cdot \dfrac{1}{\sqrt{x}}=z'\cdot\dfrac{1}{\sqrt{x}}-z\cdot\dfrac{1}{2x^{\frac{3}{2}}}=z'\cdot\dfrac{1}{\sqrt{x}}-z\cdot\dfrac{x^{-\frac{3}{2}}}{2}=z'\cdot\dfrac{1}{\sqrt{x}}+z\cdot(x^{\frac{1}{2}})'=z'\cdot\dfrac{1}{\sqrt{x}}+z\cdot(\dfrac{1}{\sqrt{x}})'=(z\cdot\dfrac{1}{\sqrt{x}})'"


Find the general solution of the linear equation:


"z=\dfrac{\int u(x)f(x)dx+C}{u(x)}=\dfrac{\int \dfrac{1}{\sqrt{x}}\cdot\frac{x}{2}dx+C}{\dfrac{1}{\sqrt{x}}}=\dfrac{\frac{1}{2}\int \sqrt{x}dx+C}{\dfrac{1}{\sqrt{x}}}=\sqrt{x}[\dfrac{1}{2}\cdot \dfrac{2x^{\frac{3}{2}}}{3}+C]=\dfrac{x^2}{3}+C\sqrt{x}."


Taking into account that z=y2, we obtain the following solution:



"y=\plusmn\sqrt{\dfrac{x^2}{3}+C\sqrt{x}}."

Now we determine the value of the constant C

 that matches the initial condition y(1)=1. We see that only a solution with a positive sign satisfies this condition. Hence,


"y=\sqrt{\dfrac{1^2}{3}+C\sqrt{1}}=\sqrt{\dfrac{1}{3}+C}=1."

This gives: "C=\dfrac{2}{3}."



"y=\sqrt{\dfrac{x^2}{3}+\dfrac{2\sqrt{x}}{3}}."